Use a graph to show how the addition of lump-sum tax (a tax independent of income)

influences the parameters of the income determination model. Graph the two systems

individually, using solid line of (1) and dashed line for (2).

(1) AE = C + I

C = 100 + 0.6 Y

Io = 40


(2) AE = C + I

C = 100 + 0.6 Yd

Io = 40

Yd = Y – T

T = 50


1
Expert's answer
2021-03-02T17:45:04-0500

"solution
\\
AE=C+I\\
C=100+0.6Y\\
I_O=40"


"AE=C+I\\
AE=100+0,6Y+40\\
AE=100+40+0.6Y\\
AE=140+0.6Y\\
0.6Y=-140\\
Y=\frac{-140}{6}\\
Y=-23.33"

we take y to be positive since income cannot be negative

"\frac{x}{y}:[{1,23.33}][{2,46.66}][{3,69.99}][{4,93.32}]"







quiz 2


"AE=C+I\\
C=100+0.6Yd\\
I_O=40\\
Yd=Y-T\\
T=50\\
AE=100+0.6Yd+40\\
Yd=Y-50\\
therefore;AE=140+0.6[Y-50]\\
AE=140+0.6Y-30\\
AE=140-30+0.6Y\\
AE=110+0.6Y\\
0.6Y=-110\\
Y=\frac{-110}{0.6}\\
y=-18.33"

we take it as a positive since its income